OpenStudy (loser66):

Texworks help I can't type probability on Texworks. Any tip is appreciated. When I type \(P(Y_1\leq 3/4, Y_2\leq 3/4)\), it says "error" . :(

OpenStudy (loser66):

@ParthKohli

OpenStudy (anonymous):

\[P(Y1 \leq \frac{3}{4}, Y2 \leq \frac{3}{4})\], It working fine for me in the equation. Is Texwork is different?

OpenStudy (loser66):

Let me take a snapshot to show how it says

OpenStudy (loser66):

You see, it didn't compile my code :(

OpenStudy (anonymous):

Let me check.

OpenStudy (anonymous):

I don't have texwork in my windows so installing and trying the found solution. So bear with me.

OpenStudy (loser66):

Sure. I am waiting. It is important to me. I have to write my graduate project on it at the end of this semester. So that I have to master it to have my job done. thanks for being here. :)

OpenStudy (anonymous):

Yep, found it. Try to use \le instead of \legslant and tell me what is happening. It is woking fine for me.

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Yes, It works now. thanks a lot.

OpenStudy (anonymous):

You could have try this too, it will work too. \[usepackage(amssymb)\]

OpenStudy (loser66):

Yes, I do. thanks again :)

OpenStudy (anonymous):

no problem :)